Determine Which of these function are bijection from the set of real number to itself



1- f(x)=-3x+4



2- f(x)=-3x2+7



3- f(x)=(x+1)/(x+2)

1
Expert's answer
2022-05-26T14:58:12-0400

1) "\forall x \in R\; \; \exist !y = -3x+4" and "\forall y \in R \; \exist!x= \frac{4-y}{3} \Rightarrow" bijection

2) "y=-3x^2+7" , "y(-1) = y(1) = 4 \Rightarrow" is not injective, thus not bijection

3) f(x)=(x+1)/(x+2). It is not a function from R to R because it is not defined at −2! So "f" certainly cannot be a bijective function.